Eliminate a and B from the equation Z=(x^2+a)(y^2+b)

1
Expert's answer
2022-02-01T16:36:52-0500
"z=(x^2+a)(y^2+b)=x^2y^2+ay^2+bx^2+ab"

Differentiate "z" with respect to "x"


"p=\dfrac{\partial z}{\partial x}=2xy^2+2bx"

Differentiate "z" with respect to "y"


"q=\dfrac{\partial z}{\partial y}=2x^2y+2ay"

Then


"pq=(2xy^2+2bx)(2x^2y+2ay)"

"=4xy(y+b)(x+a)=4xyz"

Therefore


"pq=4xyz"